# Quasar Pairs as Large–Scale Structure tracers

### v1

@@ 1 Introduction
Studies of the formation and evolution of structures in the Universe have made important recent advances from large galaxy surveys. Among the several new data sets, the Sloan Digital Sky Survey (SDSS) has made major contributions through the identification and redshift measurements of galaxies and quasars building homogeneous samples suitable for large-scale studies. The high luminosity of quasars make them ideal tracers of structure in the distant Universe and there have been works analysing the autocorrelation function in redshift space which allows for an estimate of the bias of these systems with respect to the mass distribution (White, 2012, and references therein). The results show the correlation function amplitude consistent with rich groups of galaxies, which goes in line with the expectation from numerical simulation models. Besides, these studies have shown a lack of luminosity dependence of the correlation function on quasar luminosity, a fact that is opposite to that found for galaxies where a strong clustering amplitude dependence on galaxy luminosity is observed (Zehavi et al., 2011). This independence of clustering strength on quasar luminosity is not surprising since quasar luminosity is associated to accretion onto the central black hole, a relatively stochastic process which would be relatively independent of the global environment of the quasar host when compared to the case of galaxies where luminosity produced by stars results a very good proxy of galaxy mass according to simulations. In this context, the exploration of the distribution of quasar systems and their association to mass may give new light on the joint formation and evolution of quasars and structure. The relatively low number density of quasars allows to consider mainly quasar pairs as suitable systems for statistical studies, with a strongly decaying number of systems with higher membresy. In this work we identify quasar pairs from the SDSS DR16 (Lyke, 2020) with suitable relative separation and radial velocity. In our study we adopt a maximum projected distance of $d_{p}\leq 2\text{Mpc}$ and relative velocities within $|\Delta V|\leq 900\text{km}/s$ as reasonable limits for the identification of physical pairs, with high probability of residing in the same halo. Our final sample of 2777 quasar pairs is suitable for statistical studies in both correlation function and Cosmic Microwave Background (CMB) convergence lensing maps. The samples selection are given in section 2, the statistical methods to calculate the cross–correlation function are presented in section 3 and the results for the correlation functions and CMB convergence lensing studies in section 6.
@@ 2 Quasar Catalog and Subsample
The Sloan Digital Sky Survey (SDSS) represents a stunning multi-spectral imaging and spectroscopic redshift survey, utilizing a 2.5-meter wide-angle optical telescope in conjunction with two distinct multi-fiber spectrographs, all situated in New Mexico. One of the most recent catalog, known as SDSS DR16, has successfully confirmed the detection of approximately 750414 quasars, achieved with a spectral resolution of $R=\frac{\lambda}{\Delta\lambda}\simeq 2000$ Lyke (2020). The catalog spans a significant portion of cosmic history, $7.50<z<$ 7.5, and covers approximately 14555 $\text{deg}^{2}$ of the sky [see Fig. (1)] . Regarding the absolute magnitude in the I band, it spans a large interval, $-30<M_{I}<-23$. The SDSS DR16 encompasses three distinct target fields: the Southern Stripe, the Northern Galactic Cap, and the Equatorial Stripe. Each of these areas has been meticulously selected to maximize the coverage of the survey and provide a comprehensive understanding of the quasar population across the sky Lyke (2020).
To enhance the detection of quasar pairs within the SDSS DR16 catalog, we constructed a sample that focuses on a more restricted redshift window of $2.2\leq z\leq 2.8$. The distribution of quasars as a function of the redshift is illustrated in Fig. (1). The analysis covers a total of 147325 quasars, which corresponds to approximately $20\%$ of the entire SDSS DR16 catalog  [^1] https://www.sdss4.org/dr16/algorithms/qso_catalog. There are several compelling reasons to take into account the aforementioned redshift window. Firstly, it corresponds to a pivotal epoch in the universe’s history, approximately 10 to 12 billion years ago, a critical period for unraveling the processes underpinning galaxy formation and the evolution of supermassive black holes. During that period, the Universe experienced substantial growth and dynamic activity, rendering it an optimal phase for investigating the initial stages of quasar formation. Secondly, quasars within that specific redshift range are frequently among the most luminous objects observable. Their inherent brightness facilitates detailed spectroscopic studies Richards (2006), allowing for high-quality observational data that can be leveraged to examine their physical characteristics and the environments in which they reside. Furthermore, the aforementioned range is crucial for exploring the co-evolution of galaxies and their central supermassive black holes Fanidakis (2013). Investigating the interactions between quasars and their host galaxies provides valuable insight into the mechanisms that govern galaxy growth and evolution Ross (2013). The abundance of quasars within this redshift window allows for extensive statistical analyses and comparisons, essential for deriving robust conclusions regarding the properties of quasar pairs and their significance in the broader context of cosmic history Porciani (2004).
Another important reason for our selection refers to the completeness of the sample Schneider (2003), Pâris (2012), Lyke (2020). To be more precise, there is a robust selection criterion based on brightness/magnitude and color and multi-color imaging systems that allows for effective identification of quasars by differentiating them from stars and galaxies. Second, the deep-scan capabilities help to reach a specific magnitude threshold, which ensures that quasars at higher redshifts are still detectable. Third, a spectroscopic follow-up approach is conducted after initial identification through photometric observations, confirming the quasar nature and redshift measurements. The survey covers a larger volume, and the likelihood of finding quasars increases, contributing to the completeness of the sample. All these elements and analyses help us to understand the completeness and any potential selection effects that could influence the observed distribution of quasars Schneider (2003), Pâris (2012). A typical completeness value for quasars in the SDSS DR16 catalog is often around $99.8\%$, accompanied by a contamination rate of less than $1.3\%$ Lyke (2020).
@@
Figure 1: Top panel: A projection of the 3D map representation of the quasars from the SDSS DR16Q catalog, where the color bar indicates the redshifts of each quasar. Bottom panel: We display the redshift distribution of the quasars, normalized to the total number of objects in the sample. The median value of this distribution is marked with a distinct point for clarity. The subsample distribution within the redshift range $z\in[2.2,2.8]$ is highlighted with vertical lines.
@@
The gravitational processes that dictate the formation and evolution of cosmic structures are elegantly encoded in the correlation function of galaxies. Moreover, the correlation function of quasar distributions plays a crucial role, since quasars serve as exceptional tracers of the Large–scale structures, particularly at high redshifts ($z>1$) Shen (2009), Laurent (2017). Their unique properties allow for an exploration of the gravitational dynamics that govern the evolution of the Universe, providing valuable insights that complement analyses of galaxy clustering Song (2016).
We start our investigation with an exploratory analysis of the data drawn from the SDSS DR16 catalog Lyke (2020). In this endeavor, we apply a series of cuts based on the absolute magnitude in the I-band across the sample. Data are represented according to the relationship $|M|_{I}(z)=|M|_{0}+1.8(z-0.8)$, where $|M|_{0}\in[|M|_{\text{min}},|M|_{\text{max}}]$. Here, the lower and upper limits of each band are defined as $|M|_{\text{min}}=18+n$, $|M|_{\text{max}}=|M|_{\text{min}}+1$, where $1\leq n\leq 10$ (see Fig. 2). This methodology enables us to discern any gradients present in the quasar data, which is crucial for the subsequent construction of the (weighted) random data set.
@@
Now, we proceed with the numerical computation of the two-point correlation function (2PCF), using the capabilities of the corrfunc package Sinha (2020). The corrfunc package [^2] https://github.com/manodeep/Corrfunc. allows us to understand the intricacies of the correlation function, specifically by examining the autocorrelation properties and the coherence length associated with our selected subsample. By conducting this analysis, our aim is to elucidate the power-law behavior that emerges within the linear regime White (2012). To facilitate its exploration, we initiate the process by constructing a random catalog that mirrors the redshift interval of our original data set. The random catalog serves as a crucial reference point, enabling us to compare the observed correlations with those expected in a uniformly distributed sample.
@@
For a resolution of $N_{\text{side}}=32$, we construct a mask to accurately mimic the distribution of quasars across the celestial sphere. To achieve our goal, we use the Healpix package [^3] https://healpix.sourceforge.io Górski et al (2005), where a $N_{\text{side}}=32$ is represented by 12288 pixels. The procedure is as follows: We construct a smoothed and weighted map designed for quasars to populate the random catalog, which accounts for the gradient in the data Lyke (2020). The latter smoothed map serves as a probability density function (PDF). The process begins by generating random right ascension (RA) and declination (DEC) coordinates. We then check the compatibility of these random coordinates with the smoothed map. We also ensure that each point falls within the designated mask; thus, points are more likely to be accepted into the random catalog if they are generated in regions with a high density of quasars. We evaluated the probability of the normalized weighted map at these coordinates and compared the probability with a randomly generated filter value between 0 and 1. If the filter value exceeds the probability at the selected RA/DEC coordinate, the process is restarted from the previous step. Otherwise, we accept and save the random RA and DEC. The key point is now that we are working directly with a map of probabilities instead of computing a PDF. In doing so, we ensure that the random RA/DEC points are concentrated in areas where quasars are more prevalent, resulting in a catalog that accurately reflects the spatial distribution of the original dataset. In Fig. (3), we present the distribution of quasars across the celestial sphere alongside its corresponding mask based on the procedure mentioned earlier. We also verified that the distribution of the random data in redshift mirrors the trends observed in the original data sample. The comparison ensures that our randomization process preserves the underlying characteristics of the redshift distribution, allowing us to confidently assess the significance of our findings.
To compute the correlation function in the redshift space using the corrfunc package Sinha (2020), we make use of the Landy & Szalay (1993) estimator provided it reaches minimal variance,
@@
[…]
We applied the Bootstrap method to estimate errors directly from the calculated correlation function Norberg (2009), Mohammad (2021). Specifically, we divide the survey into $N_{s}$ subsamples and perform a random resampling with replacement to create bootstrap samples $N_{s}$. For each bootstrap sample, we compute the correlation function based on the resampled data. The covariance matrix $C_{ij}$ between two estimates $\xi_{i}$ and $\xi_{j}$ at different bin values $s_{i}$ and $s_{j}$ is given by Norberg (2009), Mohammad (2021):
@@ 4 Quasar Pair Catalog
We outline the primary criteria used to identify quasar pairs within the subsample of quasars at redshifts $2.2\leq z\leq 2.8$ extracted from the SDSS DR16 catalog Lyke (2020). The process of identifying these quasar pairs through spectroscopic measurements requires a series of physically reasonable assumptions. To be consistent with the limitation given by the SDSS DR16 sample [^4] The quasar sample from SDSS DR16Q exhibits characteristic statistical redshift uncertainties on the order of $|\Delta z|\simeq 0.001$ Lyke (2020)., we propose that a pair of quasars must have the following features:
@@
  • The quasar pairs must adhere to a limiting velocity constraint such that $|\Delta V|=c(z_{2}-z_{1})\leq 900\text{km}/s$.
@@
  • The quasars that make up the pairs should be in close proximity, specifically with a projected distance of $d_{p}\leq 2\text{Mpc}$.
@@
These two conditions together define the cylindrical volume within which the quasar pairs are located. In addition, it is expected that within any given pair of quasars one will exhibit a higher luminosity, reflected by a smaller (and thus more negative) value of apparent magnitude, $m_{i}$. The method developed to identify the aforementioned quasar pairs can be summarized as follows. Each quasar in the catalog is assigned a unique ID number along with all its physical properties, including angular coordinates (RA, DEC), redshift, apparent magnitude, and luminosity expressed in solar units. We use the Planck 2018 background as a fiducial cosmology for numerical computations with the Astropy package [^5] https://www.astropy.org/ Robitaille (2013). The angular diameter distance for each quasar is calculated on the basis of its redshift and the Planck 2018 cosmology. The subtended angle in radians is determined from a specified maximum distance and is subsequently converted to degrees. The RA and DEC coordinates are translated into Healpix/Healpy [^6] https://healpy.readthedocs.io/en/latest/ pixel indices for spatial analysis Górski et al (2005). Each pixel corresponds to a specific location in the celestial sphere, and we employ a high resolution of $N_{\text{side}}=2048$. We store the vectors for each pixel for future use in the query-disk routine and create a list to keep track of neighboring pixels. For each quasar, neighboring pixels within a defined angular distance are identified, facilitating the analysis of local quasar environments. The radius in radians is computed using the angular distance, and the neighboring pixels within the specified disk are recorded. The algorithm then searches for nearby quasar pairs based on their redshift and angular distance, ensuring that no pair is identical and that they lie within a predetermined redshift tolerance. Filter out duplicate pairs employing ordered tuples and retains only the closest pairs, constrained by a physical distance criterion of $d_{p}\leq 2\text{Mpc}$.
Figure 5: The distribution of the computed projected distances of quasar pairs.

### v2

@@ 1 Introduction
Studies of the formation and evolution of structures in the Universe have made important recent advances from large galaxy surveys. Among the several new data sets, the Sloan Digital Sky Survey (SDSS) has made major contributions through the identification and redshift measurements of galaxies and quasars building homogeneous samples suitable for large-scale studies Lyke et al. (2020).
The high luminosity of quasars make them ideal tracers of structure in the distant Universe and there have been works analysing the autocorrelation function in redshift space which allows for an estimate of the bias of these systems with respect to the mass distribution (Porciani et al., 2004; White et al., 2012, and references therein). The results show the correlation function amplitude consistent with rich groups of galaxies, which goes in line with the expectation from numerical simulation models. Besides, these studies have shown a lack of luminosity dependence of the correlation function on quasar luminosity, a fact that is opposite to that found for galaxies where a strong clustering amplitude dependence on galaxy luminosity is observed (Zehavi et al., 2011). This independence of clustering strength on quasar luminosity is not surprising since quasar luminosity is associated to accretion onto the central black hole, a relatively stochastic process which would be relatively independent of the global environment of the quasar host when compared to the case of galaxies where luminosity produced by stars results a very good proxy of galaxy mass according to simulations Alberts &\&& Noble (2022).
In this context, the exploration of the distribution of quasar systems and their association to mass may give new light on the joint formation and evolution of quasars and structure. The relatively low number density of quasars allows to consider mainly quasar pairs as suitable systems for statistical studies, with a strongly decaying number of systems with higher membresy.
In this work we identify quasar pairs from the SDSS DR16 (Lyke et al., 2020) with suitable relative separation and radial velocity. In our study we adopt a maximum projected distance of $d_{p}\leq 2\text{Mpc}$ and relative velocities within $|\Delta V|\leq 900\text{km}/s$ as reasonable limits for the identification of physical pairs, with high probability of residing in the same halo. Our final sample of 2777 quasar pairs is suitable for statistical studies in both correlation function and Cosmic Microwave Background (CMB) convergence lensing maps. The selection criteria for the quasar samples are detailed in Section 2. The statistical methodologies employed to compute the autocorrelation of the quasar sample are outlined in Section 3. The quasar catalog, specifically for the quasar pairs, is provided in Section 4. Following this, Section 5 focuses on the characterization of luminosity and color-color maps. Section 6 presents the cross-correlation functions, while the results pertaining to the studies on CMB convergence lensing are discussed in Section 7.
@@ 2 Quasar Catalog and Subsample
The Sloan Digital Sky Survey (SDSS) represents a stunning multi-spectral imaging and spectroscopic redshift survey, utilizing a 2.5-meter wide-angle optical telescope in conjunction with two distinct multi-fiber spectrographs, all situated in New Mexico. One of the most recent catalog, known as SDSS DR16, has successfully confirmed the detection of approximately 750414 quasars, achieved with a spectral resolution of $R=\frac{\lambda}{\Delta\lambda}\simeq 2000$ Lyke et al. (2020). The catalog spans a significant portion of cosmic history, $7.50<z<$ 7.5, and covers approximately 14555 $\text{deg}^{2}$ of the sky [see Fig. (1)] . Regarding the absolute magnitude in the I band, it spans a large interval, $-30<M_{I}<-23$. The SDSS DR16 encompasses three distinct target fields: the Southern Stripe, the Northern Galactic Cap, and the Equatorial Stripe. Each of these areas has been meticulously selected to maximize the coverage of the survey and provide a comprehensive understanding of the quasar population across the sky Lyke et al. (2020).
To enhance the detection of quasar pairs within the SDSS DR16 catalog, we constructed a sample that focuses on a more restricted redshift window of $1.2\leq z\leq 2.8$. The distribution of quasars as a function of the redshift is illustrated in Fig. (1). The analysis covers a total of 147325 quasars, which corresponds to approximately $66\%$ of the entire SDSS DR16 catalog  [^1] https://www.sdss4.org/dr16/algorithms/qso_catalog. There are numerous compelling reasons to take into account the redshift window previously mentioned. We will delineate a series of justifications grounded in theoretical frameworks as well as numerical analyses derived from various catalogs utilized in the existing literature. It corresponds to a pivotal epoch in the universe’s history, approximately 10 to 12 billion years ago, a critical period for unraveling the processes underpinning galaxy formation and the evolution of supermassive black holes (SMBHs). During this period, the Universe underwent significant growth and activity, making it an ideal time to explore the early stages of quasar formation. Long ago, it was proposed that supermassive black holes (SMBHs) originate from the collapse of massive gas clouds in the early Universe. The energy released during the accretion process plays a crucial role in galaxy evolution, highlighting the importance of feedback mechanisms in regulating star formation within galaxies Silk &\&& Rees (1998).
Subsequently, the connection between black hole mass and the velocity dispersion of the host galaxy bulge-known as the $M_{BH}-\sigma$ relation-has been essential in showing that black hole growth is closely tied to the evolution of its host bulge. This relationship has been studied extensively in the context of hierarchical galaxy formation models, further illustrating the interconnectedness of these cosmic processes Haehnelt &\&& Kauffmann (2000), Merloni &\&& Heinz (2007). The coevolution of SMBHs and their host galaxies often results from significant mergers between galaxies, which produce feedback that affects both the black hole and the galaxy itself Kormendy &\&& Ho (2013), Harrison et al. (2018).
[…]
Another important aspect of quasars is how their number density changes with redshift, which is described by the quasar luminosity function (QLF). The QLF measures the number density of quasars per unit luminosity and is usually expressed as a double power-law, featuring a luminosity break, $L_{\star}$. This framework distinguishes two scenarios: in luminosity evolution, the luminosity varies over time while the number density stays constant, whereas in density evolution, the number density changes while the luminosity of individual quasars remains stable Richards et al. (2006). Measuring the QLF is crucial for understanding the evolution of supermassive black holes (SMBHs) and for placing constraints on various cosmic backgrounds Ross et al. (2013). A detailed analysis of the QLF was carried out using the SDSS-III DR9 catalog, which includes 22,301 quasars across an area of $2236\rm{deg}^{2}$. The catalog provides spectroscopic redshift measurements in the range of $z\in[2.2,3.5]$ and incorporates K-corrections for accurate luminosity assessments. The results confirmed the double power-law relationship in the range $z\in[0.3,2.2]$], but also identified a clear break in this relationship around $z=2$ with $M_{I}(z=2)=-24.5$. For $z\geq 2.2$ , they found a log-linear trend in quasar luminosity amplitude and break luminosity Ross et al. (2013).
By combining data from the SDSS and 2dFQSO surveys, researchers examined the clustering properties of quasars up to redshift $z=5$, calculating the two-point projected correlation function for quasars within $r<20\text{h}^{-1}\text{Mpc}$. Their analysis indicated a halo mass of about $10^{13}M_{\odot}$, with a slight increase observed with redshift. Additionally, when considering the fraction of halos hosting active quasars, the typical lifetime of quasars at high redshift ($z>1$) is estimated to be around $t_{Q}\simeq 10^{8}\rm{yr}$ Porciani et al. (2004).
[…]
@@
Figure 1: Top panel: A projection of the 3D map representation of the quasars from the SDSS DR16Q catalog, where the color bar indicates the redshifts of each quasar. Bottom panel: We display the redshift distribution of the quasars, normalized to the total number of objects in the sample. The median value of this distribution is marked with a distinct point for clarity. The subsample distribution within the redshift range $z\in[1.2,2.8]$ is highlighted with vertical lines.
@@
The gravitational processes that dictate the formation and evolution of cosmic structures are elegantly encoded in the correlation function of galaxies. Moreover, the correlation function of quasar distributions plays a crucial role, since quasars serve as exceptional tracers of the Large–scale structures, particularly at high redshifts ($z>1$) Shen et al. (2009), Laurent et al. (2017). Their unique properties allow for an exploration of the gravitational dynamics that govern the evolution of the Universe, providing valuable insights that complement analyses of galaxy clustering Song et al. (2016).
We start our investigation with an exploratory analysis of the data drawn from the SDSS DR16 catalog Lyke et al. (2020). In this endeavor, we apply a series of cuts based on the absolute magnitude in the I-band across the sample. Data are represented according to the relationship $|M|_{I}(z)=|M|_{0}+1.8(z-0.8)$, where $|M|_{0}\in[|M|_{\text{min}},|M|_{\text{max}}]$. Here, the lower and upper limits of each band are defined as $|M|_{\text{min}}=18+n$, $|M|_{\text{max}}=|M|_{\text{min}}+1$, where $1\leq n\leq 10$ (see Fig. 2). This methodology enables us to discern any gradients present in the quasar data, which is crucial for the subsequent construction of the (weighted) random data set.
@@
Now, we proceed with the numerical computation of the two-point correlation function (2PCF), using the capabilities of the corrfunc package Sinha & Garrison. (2020). The corrfunc package [^2] https://github.com/manodeep/Corrfunc. allows us to understand the intricacies of the correlation function, specifically by examining the autocorrelation properties and the coherence length associated with our selected subsample. By conducting this analysis, our aim is to elucidate the power-law behavior that emerges within the linear regime White et al. (2012). To facilitate its exploration, we initiate the process by constructing a random catalog that mirrors the redshift interval of our original data set. The random catalog serves as a crucial reference point, enabling us to compare the observed correlations with those expected in a uniformly distributed sample.
@@
For a resolution of $N_{\text{side}}=32$, we construct a mask to accurately mimic the distribution of quasars across the celestial sphere. To achieve our goal, we use the Healpix package [^3] https://healpix.sourceforge.io Górski et al. (2005), where a $N_{\text{side}}=32$ is represented by 12288 pixels. The procedure is as follows: We construct a smoothed and weighted map designed for quasars to populate the random catalog, which accounts for the gradient in the data Lyke et al. (2020). The latter smoothed map serves as a probability density function (PDF). The process begins by generating random right ascension (RA) and declination (DEC) coordinates. We then check the compatibility of these random coordinates with the smoothed map. We also ensure that each point falls within the designated mask; thus, points are more likely to be accepted into the random catalog if they are generated in regions with a high density of quasars. We evaluated the probability of the normalized weighted map at these coordinates and compared the probability with a randomly generated filter value between 0 and 1. If the filter value exceeds the probability at the selected RA/DEC coordinate, the process is restarted from the previous step. Otherwise, we accept and save the random RA and DEC. The key point is now that we are working directly with a map of probabilities instead of computing a PDF. In doing so, we ensure that the random RA/DEC points are concentrated in areas where quasars are more prevalent, resulting in a catalog that accurately reflects the spatial distribution of the original dataset. In Fig. (3), we present the distribution of quasars across the celestial sphere alongside its corresponding mask based on the procedure mentioned earlier. We also verified that the distribution of the random data in redshift mirrors the trends observed in the original data sample. The comparison ensures that our randomization process preserves the underlying characteristics of the redshift distribution, allowing us to confidently assess the significance of our findings.
To compute the correlation function in the redshift space using the corrfunc package Sinha & Garrison. (2020), we make use of the Landy & Szalay (1993) estimator provided it reaches minimal variance,
@@
Figure 4: Top panel: Two-point correlation function obtained using the LS estimator for the total sample, $z\in[1.2,2.8]$. Red point represents the correlation function with its error bars. Bottom panel: The dashed-red curve indicate a best linear fit in the logarithm scale. In the linear regime, the correlation function can be parametrized as $\xi(s)=(s/s_{0})^{-\gamma}$ with $\gamma=1.21\pm 0.03$ and $s_{0}=12.25\pm 1.30$.
We applied the Bootstrap method to estimate errors directly from the calculated correlation function Norberg et al. (2009), Mohammad & Percival (2021). Specifically, we divide the survey into $N_{s}$ subsamples and perform a random resampling with replacement to create bootstrap samples $N_{s}$. For each bootstrap sample, we compute the correlation function based on the resampled data. The covariance matrix $C_{ij}$ between two estimates $\xi_{i}$ and $\xi_{j}$ at different bin values $s_{i}$ and $s_{j}$ is given by Norberg et al. (2009), Mohammad & Percival (2021):
@@ 4 Quasar Pair Catalog
We outline the primary criteria used to identify quasar pairs within the subsample of quasars at redshifts $1.2\leq z\leq 2.8$ extracted from the SDSS DR16 catalog Lyke et al. (2020). The process of identifying these quasar pairs through spectroscopic measurements requires a series of physically reasonable assumptions. To be consistent with the limitation given by the SDSS DR16 sample [^4] The quasar sample from SDSS DR16Q exhibits characteristic statistical redshift uncertainties on the order of $|\Delta z|\simeq 0.001$ Lyke et al. (2020)., we propose that a pair of quasars must have the following features:
@@
  • The quasar pairs must adhere to a limiting velocity constraint such that $|\Delta V|=c(z_{2}-z_{1})\leq 900\rm{km}/s$.
@@
  • The quasars that make up the pairs should be in close proximity, specifically with a projected distance of $d_{p}\leq 2\rm{Mpc}$.
@@
These two conditions together define the cylindrical volume within which the quasar pairs are located. In addition, it is expected that within any given pair of quasars one will exhibit a higher luminosity, reflected by a smaller (and thus more negative) value of apparent magnitude, $m_{i}$. The method developed to identify the aforementioned quasar pairs can be summarized as follows. Each quasar in the catalog is assigned a unique ID number along with all its physical properties, including angular coordinates (RA, DEC), redshift, apparent magnitude, and luminosity expressed in solar units. We use the Planck 2018 background as a fiducial cosmology for numerical computations with the Astropy package [^5] https://www.astropy.org/ Robitaille et al. (2013). The angular diameter distance for each quasar is calculated on the basis of its redshift and the Planck 2018 cosmology. The subtended angle in radians is determined from a specified maximum distance and is subsequently converted to degrees. The RA and DEC coordinates are translated into Healpix/Healpy [^6] https://healpy.readthedocs.io/en/latest/ pixel indices for spatial analysis Górski et al. (2005). Each pixel corresponds to a specific location in the celestial sphere, and we employ a high resolution of $N_{\text{side}}=2048$. We store the vectors for each pixel for future use in the query-disk routine and create a list to keep track of neighboring pixels. For each quasar, neighboring pixels within a defined angular distance are identified, facilitating the analysis of local quasar environments. The radius in radians is computed using the angular distance, and the neighboring pixels within the specified disk are recorded. The algorithm then searches for nearby quasar pairs based on their redshift and angular distance, ensuring that no pair is identical and that they lie within a predetermined redshift tolerance. Filter out duplicate pairs employing ordered tuples and retains only the closest pairs, constrained by a physical distance criterion of $d_{p}\leq 2\text{Mpc}$.
Figure 5: Top panel: Distribution of the number of quasar pairs across the complete redshift window of the SDSS DR16 quasar catalog. The highest concentration of pairs is found within the redshift range $z\in[1.2,2.8]$, totaling 2777 pairs. Overall, the catalog contains a maximum of 4011 pairs. Bottom panel: The distribution of quasar pairs detected within a cylindrical volume characterized by a projected radius of $d_{p}\leq 2\text{Mpc}$ and a length of $|\Delta V|=900\text{km}/s$ is presented. The black vertical lines denotes $d^{\text{min}}_{p}\simeq 0.03\text{Mpc}$, representing the minimum projected separation at which a quasar pair can be reliably identified. This threshold is determined by the angular resolution of the SDSS camera, set at $\theta_{SDSS}=5^{"}$ along with a mean redshift value of $z_{\text{mean}}=2.4$.
The distribution of the identified quasar pairs is shown in Fig. (5). In the SDSS DR16 catalog, with redshifts ranging from $z=0.0007$ to $z=7.023$, a total of 4011 quasar pairs have been identified. Most of these quasar pairs are found in the first two redshift bins: $z\in[0.0007,1.2]$ and $z\in[1.2,2.8]$ with pairs detected beyond $z=4.4$. The highest detection rate occurs in the second redshift interval $z\in[1.2,2.8]$, accounting for a sample of 2777 pairs where we will focus our study. For example, the histogram of projected distances in Mpc for all 2777 quasar pairs is also shown in Fig. (5). The average distance between pairs is $d=(0.986\pm 0.5009)\text{ Mpc}$ with only $21$ pairs have projected separations below the minimum projected distance resolved by the SDSS, $d^{\text{min}}_{p}\simeq 0.03\text{Mpc}$. Overall, the peak of the histogram suggests that quasar pairs tend to cluster around this average projected distance.